for a particular event 784 tickets were sold for a total of $4052. If students paid $4 perticket and nonstudents paid $6 per ticket, how many student tickets were sold?
Answer by addingup(3677)

student: x
non student: y
x+y = 784 subtract x from both sides:
y = 784-x
--------------------
4x+6y = 4052 substitute for y:
4x+6(784-x) = 4052
4x+4704-6x = 4052
-2x = -652
x = 326 student tickets were sold.